OpenStudy (lena772):

The regular price of a child's entry ticket to a water park is $8 less than that for an adult's. The park offers half off of all entry tickets during the off-peak season. The Sandlers paid a total of $194 for 1 adult ticket and 4 child's tickets to the water park during the off-peak season. The following equation represents this situation, where x represents the regular price of an adult ticket. 194 = one-halfx + 2(x - 8) What is the regular price of a child's ticket? $68 $76 $84 $89

OpenStudy (anonymous):

Adult ticket = x Child ticket = x-8 if 194=0.5x+2(x-8) Total = 1 adult (half price) + 4 kids(half price ie 2 kids) they should have paid (peak time) 388 = x + 4(x-8) 388 = x + 4x -32 420=5x x=84 (adult) so child is x-8 = 84-8=?

OpenStudy (lena772):

76

OpenStudy (anonymous):

yep!